Answer both of the following Roman numeral questions (I, II). Equal credit. A company has fixed cost of $44,000, variable cost per unit of S12, and sells its product at S19 each a. What quantity must the firm sell in order to break even? Explain how you reached this conclusion. b. What quantity must the firm sell in order to make a profit of $43,000? Explain how you reached this conclusion Hello I need help with this question. Please answer both questions and please show all work so I have a better understanding how to better do this. Thanks please show all work.
Answer #1

a) Break even quanity = Fixed Cost/[ Selling price- Cost price] = 44000/[19-12]=44000/7= 6285.7

6286 must be sold

b) Profit+Fixed cost = 44000+43000=87000

Quanity = 87000/7=12428.57